What is the electric potential at point P?
Answer #1

DATA:

q_{1}=-6\, nC=-6\times 10^{-9}\, C

q_{2}=+2\, nC=+2\times 10^{-9}\, C

r_{1}=r_{2}=r=1\, cm=1\times 10^{-2}\: m

SOLUTION:

The electric potential is defined as:

V=k\frac{q}{r}

In this problem, for a distribution of charges, the total potential at point P , V_{p}, is defined:

V_{p}=k\frac{q_{1}}{r_{1}}+k\frac{q_{2}}{r_{2}}

or

V_{p}=\frac{k}{r}\left ( q_{1}+q_{2} \right )

Replacing values given in DATA:

V_{p}=\frac{9\times 10^{9}\, N.m^{2}/C^{2}}{1\times 10^{-2}m}\left ( -6\times 10^{-9}\, C+2\times 10^{-9}\, C \right )

{\color{Blue}V_{p}=-3600\, V \approx -3.6\times 10^{3}\, V}